DETERMINE THE NUMBER OF SETS OF ALL POSSIBLE RANDOM SAMPLES THAT CAN BE DRAWN FROM THE GIVEN POPULATION BY USING THE FORMULA, NCn.



N = 10, n = 3

1
Expert's answer
2022-05-12T07:32:12-0400

The number of possible samples which can be selected without replacement is

"_NC_n=\cfrac{N! } {n! \cdot(N-n)! };"

"_{10}C_3=\cfrac{10! } {3! \cdot(10-3)! }=\cfrac{8\cdot9\cdot10}{2\cdot3}=120."